# Bayesian Projection of Refugee and Asylum Seeker Populations

### v1

@@ Methods
where $d_{\mathrm{Exp}}$ is the density of the Exponential distribution and $d_{N}$ the density of the normal distribution. A final peak position $\delta^{*}$ is drawn with probabilities $p_{\delta}$, and the final projections are set to $\mu^{(k)}_{c[m],t}=\mu_{c[m],t,\delta^{*}}^{(k)}$. Figure 3 illustrates how projections for ongoing growth crises are generated conditional on $\delta$.
@@
Figure 3: Diagram of projections for ongoing crisis in the growth phase. Projections are generated conditional on a parameter $\delta$ that controls how long the crisis will continue before reaching its peak.
@@ Results
As a validation exercise, we held out all observations after a cutoff time point $t^{*}$ and applied the full projection pipeline. The validation set included all countries that did not have a new crisis start after the cutoff time point.
The mean error and mean absolute error for the proposed Bayesian method and the benchmark deterministic logistic method were computed for $1$-year, $5$-year, and $10$-year projection horizons. The results are presented in Table 2. For the $1$-year ahead projections, the Bayesian method had lower mean error and mean absolute error than the benchmark. The benchmark performed slightly better in terms of mean error for the $5$- and $10$-year projections than the Bayesian method.
@@
Projections for all countries with an ongoing crisis detected as of 2021 are included in the appendix. Illustrative projections for six of these countries are shown in Figure 4. Of these countries, Colombia, Gambia, and Somalia have crises in the decline phase, and El Salvador, Haiti, and Honduras have crisis in the growth phase. The strength of the probablistic projections from the Bayesian model as compared to the deterministic benchmark can be seen in Honduras. Despite the posterior median projection being similar to the deterministic projection, the Bayesian credible intervals indicate the possibility of continued growth in the refugee and asylum seeker population. In Gambia, the posterior median projects faster decline than the deterministic benchmark because the Bayesian model incorporates the previous rates of decline observed in the crisis.
@@
Figure 5 illustrates how projections for El Salvador would have changed year over year as new data became available. In 2016, the 95% credible intervals include the possibility of an increase in the refugee population to around the same level as the peak of the previous crisis. For the 2017 estimates, the model adjusts upwards its projections based on the increased rate of change implied by the new data point. By 2021, the posterior median suggests the crisis has reached its peak, although there is still significant posterior mass on the possibility of continued increases up a level last seen in the first refugee crisis in the country in the 1980s.
@@ Discussion
We have proposed a Bayesian hierarchical time-series model of refugee and asylum seeker populations by country of origin. We then used the model to project future combined refugee and asylum seeker populations for countries experiencing ongoing refugee crises. Validation exercises suggest the model is well-calibrated at multiple time horizons.
@@
Current UN population projections assume that two thirds of refugees will return to their country of origin within 5 years (United Nations, 2022a). Our validation results show that this rule of thumb performs well for 5-year ahead projections. The Bayesian modeling approach achieves comparable or better performance in all the validation exercises, and has the additional benefit of providing well-calibrated probabilistic projections.
@@
A basic assumption of any statistical forecasting method is that the future will be in some way similar to the past. In this work, this assumption manifests itself in the model specification that assumes current refugee crises will follow a similar shape to previous crises. However, the underlying dynamics of forced displacement may change in the future in ways that fundamentally change refugee population trends. Global climate change is a prominent example of a factor that may drive future displacement in ways that differ from past causes of displacement (Piguet et al., , 2011). Uncertainty estimates from our modeling approach, however, are conditional on the model specification and do not incorporate uncertainty in how displacement dynamics may change in the future. As such, while credible intervals from our approach appear to be relatively well-calibrated in the projection horizons considered in the validation exercises, there is no guarantee that the model specification will lead to good performance in the long term.
@@
Our work leaves open multiple directions for future research. We did not attempt to forecast the outbreak of new crises, rather focusing on projecting the end of already existing crises. This focus encodes an optimistic assumption that, in the long term, all refugee crises will end. However, research has been carried out on identifying where and when future refugee crises will occur (OECD, , 2019). Producing accurate “early warnings” of crises is notoriously difficult due to the complexity of causes of refugee crises (Schmeidl and Jenkins, , 1996). An area for future research is to expand our approach to incorporate uncertainty about whether new crises will arise in the future. In addition, incorporating estimates of refugee and asylum seeker populations in country-level population projections requires projecting destination countries. We plan to address this in the future, drawing on developments in projecting bilateral migration flows between countries (Welch and Raftery, , 2022).
@@ Acknowledgements:
This research was supported by NIH grant R01 HD070936. The authors thank Patrick Gerland for helpful conversations.
@@
Hyperpriors
Peak values:
$\rho$ $\sim N(0,3),$ (18)
$\sigma_{\rho}$ $\sim\mathrm{InvGamma}(0.1,0.1).$ (19)
Peak duration:
$\psi$ $\sim\mathrm{InvGamma}(0.1,0.1).$ (21)
Growth and decline rates:
$\mu_{\omega^{(1)}}$ $\sim N(0,3),$ (22)
$\mu_{\omega^{(2)}}$ $\sim N(0,3),$ (23)
$\sigma_{\omega^{(1)}}$ $\sim\mathrm{InvGamma}(0.1,0.1),$ (24)
$\sigma_{\omega^{(2)}}$ $\sim\mathrm{InvGamma}(0.1,0.1).$ (25)
Growth and decline noise:
$\mu_{\sigma^{(1)}}$ $\sim N(-7,4),$ (26)
$\mu_{\sigma^{(2)}}$ $\sim N(-7,4),$ (27)
$\sigma_{\sigma^{(1)}}$ $\sim\mathrm{InvGamma}(0.1,0.1)$ (28)
$\sigma_{\sigma^{(2)}}$ $\sim\mathrm{InvGamma}(0.1,0.1).$ (29)
Observed crisis lengths
Figure 6: Empirical distribution of the length of the growth phase for all crises classified as ended.
Ongoing crisis predictions
The plots below show the posterior predictive distributions for the refugee proportions $\mu_{m,t}$ for all crises and time points. To simplify presentation, multiple crises from the same country are shown in the same plot. See pages 1- of plots/ongoing_crises_2021.pdf
0.1 Crisis Segmentation

### v2

@@ Methods
where $f_{\mathrm{Exp}}$ is the density of the Exponential distribution and $f_{N}$ the density of the normal distribution. A final peak position $\delta^{*}$ is drawn with probabilities $p_{\delta}$, and the final projections are set to $\mu_{c[m],t,k}=\mu_{c[m],t,\delta^{*},k}$. Figure 3 illustrates how projections for ongoing growth episodes are generated conditional on $\delta$.
@@
Figure 3: Diagram of projections for ongoing episode in the growth phase. Projections are generated conditional on a parameter $\delta$ that controls how long the episode will continue before reaching its peak.
@@ Results
As a validation exercise, we held out all observations after a cutoff time point $t^{*}$ and applied the full projection pipeline. The validation set included all countries that did not have a new episode start after the cutoff time point.
The mean error and mean absolute error for the proposed Bayesian method and the benchmark deterministic logistic method were computed for $1$-year, 55-year, and $10$-year projection horizons. The results are presented in Table 2. For the $1$-year ahead projections, the Bayesian method had lower mean error and mean absolute error than the benchmark. The benchmark performed slightly better in terms of mean error for the $5$- and $10$-year projections than the Bayesian method.
@@
Projections for all countries with an ongoing episode detected as of 2021 are included in the appendix. Illustrative projections for six of these countries are shown in Figure 4. Of these countries, Colombia, Gambia, and Somalia have episode in the decline phase, and El Salvador, Haiti, and Honduras have episodes in the growth phase. The strength of the probablistic projections from the Bayesian model as compared to the deterministic benchmark can be seen in Honduras. Despite the posterior median projection being similar to the deterministic projection, the Bayesian credible intervals indicate the possibility of continued growth in the refugee and asylum seeker population. In Gambia, the posterior median projects faster decline than the deterministic benchmark because the Bayesian model incorporates the previous rates of decline observed in the episode.
@@
Figure 5 illustrates how projections for El Salvador would have changed year over year as new data became available. In 2016, the 95% credible intervals include the possibility of an increase in the refugee and asylum seeker population to around the same level as the peak of the previous episode. For the 2017 estimates, the model adjusts upwards its projections based on the increased rate of change implied by the new data point. By 2021, the posterior median suggests the episode has reached its peak, although there is still significant posterior mass on the possibility of continued increases up a level last seen in the first episode in the country in the 1980s.
@@ Discussion
We have proposed a Bayesian hierarchical time-series model of refugee and asylum seeker population (stocks) by country of origin based on UNHCR data. We then used the model to project future refugee and asylum seeker population statistics for countries of origin of large refugee and asylum seeker populations. Validation exercises suggest the model is well-calibrated at multiple time horizons.
@@
UN population projections assume that two thirds of refugees will return to their country of origin within 5 years (United Nations, 2022a). Our validation results show that this rule of thumb performs well for 5-year ahead projections. The Bayesian modeling approach achieves comparable or better performance in all the validation exercises, and has the additional benefit of providing well-calibrated probabilistic projections.
A fundamental limitation of our projections is that they are generated based on official statistics gathered and disseminated by the UNHCR, using their definition of refugees and asylum seekers and based on their data sources and data collection methods. However, constructivist legal definitions of refugee and asylum seeker categories exclude populations who may fall into those same categories in a broader sociological sense. Furthermore, adjudication of refugee and asylum seeker status are inherently political (FitzGerald and Arar, , 2018). The UNHCR may, for example, end recognition of refugee status based on a determination of improved circumstances in a country or origin (Fitzpatrick and Bonoan, , 2003). Affected individuals would no longer be counted in official statistics, even if they would still be considered a refugee under alternative definitions. The inherent limitations of the data imply that our projections can only be interpreted as projections of refugee and asylum seeker official statistics, rather than projections of refugee and asylum seeker populations in a broader sense. While our modeling approach is likely applicable to the latter task, generating projections would require alternative data sources.
@@
A basic assumption of any statistical forecasting method is that the future will be in some way similar to the past. In this work, this assumption manifests itself in the model specification that assumes current growth and decline periods will follow a similar shape to previous episodes. However, the underlying dynamics driving refugee and asylums seeker dynamics may change in the future in ways that fundamentally change refugee population trends. Global climate change is a prominent example of a factor that may drive future displacement in ways that differ from past causes of displacement (Piguet et al., , 2011). Uncertainty estimates from our modeling approach, however, are conditional on the model specification and do not incorporate uncertainty in how displacement dynamics may change in the future. As such, while credible intervals from our approach appear to be relatively well-calibrated in the projection horizons considered in the validation exercises, there is no guarantee that the model specification will lead to good performance in the long term.
@@
Our work leaves open multiple directions for future research. We focused on projecting combined refugee and asylum seeker populations based on definitions adopted by the UNHCR and reflected in official statistics; it may be of interest to examine alternative subsets of forced migrants and using other data sources. Our modeling method is adaptable to consider other populations that follow dynamics captured by our conceptual model. An important limitation of our work is that it focuses only on forecasting how current refugee and asylum seeker counts will evolve, and, hopefully, end. We did not attempt to forecast the start of large new refugee or asylum seeker populations, rather focusing on projecting how existing large populations will change over time. This focus encodes an optimistic assumption that, in the long term, all refugee and asylum seeker populations will go to zero by those affected returning home or losing status in other ways, such as naturalisation. However, research has been carried out on identifying where and when future episodes will occur (OECD, , 2019). Producing accurate “early warnings” is notoriously difficult due to the complexity of causes of refugee migration (Schmeidl and Jenkins, , 1996), although progress has been made (Napierała et al., , 2022). An area for future research is to expand our approach to incorporate uncertainty about whether new crises will arise in the future. In addition, incorporating estimates of refugee and asylum seeker statistics in country-level population projections requires forecasting in which destination countries refugee and asylum seekers will seek protection. To address this in the future, it could be useful to draw on developments in projecting bilateral migration flows between countries (Welch and Raftery, , 2022).
@@ Acknowledgements:
This research was supported by NIH grant R01 HD070936. The authors thank Patrick Gerland for helpful conversations. Both authors thanks the Laboratoire MAP5 at Université Paris-Cité for warm hospitality.
@@
Observed episode lengths
Figure A1: Empirical distribution of the length of the growth phase for all episodes classified as ended.
Sensitivity analysis
We conducted a sensitivity analysis on the threshold for defining an episode. In the main text we consider a threshold of 0.025%. As a sensitivity analysis, we conducted the complete analysis using thresholds of 0.0125% and 0.05%.
At a threshold of 0.0125%, 9 additional episodes were detected compared to using the 0.025% threshold. At a threshold of 0.05%, 7 fewer episodes were detected compared to using the 0.025% threshold. Table 4 shows mean error and mean absolute error for the Bayesian and benchmark logistic method for both thresholds, and Table 5 shows empirical coverage results. The Bayesian method typically had lower MAE and ME for both alternative thresholds compared to the benchmark method, particularly in the 1-year ahead validations.
MAE ME
Cutoff Target $n$ Bayes Logistic Bayes Logistic
Threshold 0.00125%
1 year ahead
2016 2017 33 0.55 0.64 0.47 0.59
2017 2018 35 0.15 0.21 0.03 0.18
2018 2019 35 0.15 0.19 0.03 0.18
2019 2020 36 0.14 0.24 0.04 0.20
2020 2021 37 0.23 0.35 0.15 0.30
5 year ahead
2011 2016 19 1.25 1.35 0.76 0.68
2016 2021 30 1.40 1.34 1.25 1.31
10 year ahead
2011 2021 16 1.87 1.89 1.56 1.48
Threshold 0.5%
1 year ahead
2016 2017 23 0.77 0.88 0.65 0.80
2017 2018 23 0.23 0.31 0.06 0.27
2018 2019 24 0.19 0.27 0.03 0.25
2019 2020 25 0.17 0.30 0.02 0.25
2020 2021 27 0.30 0.47 0.20 0.40
5 year ahead
2011 2016 12 1.93 2.09 1.16 1.06
2016 2021 20 2.01 1.91 1.81 1.85
10 year ahead
2011 2021 10 2.91 2.96 2.43 2.31
Table 5: Empirical coverage of 80%, 90%, and 95% credible intervals for projections in the validation exercises for additional values of the episode threshold.
Coverage
Cutoff Target $n$ 80% 90% 95%
Threshold 0.025%
1 year ahead
2016 2017 33 87.9% 90.9% 90.9%
2017 2018 35 100.0% 100.0% 100.0%
2018 2019 35 94.3% 97.1% 100.0%
2019 2020 36 94.4% 94.4% 97.2%
2020 2021 37 86.5% 94.6% 97.3%
Average 92.6% 95.5% 97.2%
5 year ahead
2011 2016 19 73.7% 78.9% 89.5%
2016 2021 30 83.3% 90.0% 96.7%
10 year ahead
2011 2021 16 62.5% 68.8% 75.0%
Threshold 0.05%
1 year ahead
2016 2017 23 87.0% 87.0% 87.0%
2017 2018 23 100.0% 100.0% 100.0%
2018 2019 24 95.8% 95.8% 100.0%
2019 2020 25 96.0% 96.0% 100.0%
2020 2021 27 85.2% 85.2% 88.9%
Average 92.6% 92.6% 95.1%
5 year ahead
2011 2016 12 66.7% 75.0% 83.3%
2016 2021 20 85.0% 85.0% 100.0%
10 year ahead
2011 2021 10 70.0% 80.0% 80.0%
Ongoing episode predictions
Figure A2: Posterior predictive distributions for the proportions $\mu_{m,t}$ for all episodes and time points. To simplify presentation, multiple episodes from the same country are shown in the same plot.
All detected episodes
Figure A3: Detected episodes in each country according to the deterministic rules described in the main text (using the the threshold of 0.025% as in the primary analysis). Each episode is further annotated with the detected growth and (when applicable) decline phases and peaks.